Please provide excel step by step instructionsQuestion Help * 9.3.21 The number of initial public offerings of stock issued in a 10-year period and the total proceeds of t
Answer #1

X Value=   624  
Confidence Level=   95%  
      
      
Sample Size , n=   10  
Degrees of Freedom,df=n-2 =   8  
critical t Value=tα/2 =   2.306   [excel function: =t.inv.2t(α/2,df) ]

X (x-x̅)²
402 4583.29
450 13386.49
687 124397.29
480 21228.49
478 20649.69
381 2180.89
58 76341.69
57 76895.29
191 20534.89
159 30730.09

here, x̅ = ΣX/n = 3343/10 =   334.30

SSxx =    Σ(x-x̅)² =    390928.10

Standard Error of the Estimate,Se=   6933.50

margin of error,E=t*S(ŷ)= t*Se*√ (1+1/n+(X-X̅)²/Sxx) = 18332.5837
predicted value of Y at X = 324 is

Ŷ =   16963.740 +   34.962 *324 = 38780.333


Prediction Interval Lower Limit=Ŷ -E =   20447.749023 millions
Prediction Interval Upper Limit=Ŷ +E =   57112.916368 millions

so, prediction interval is ($20447749023, $57112916368)